n^2/3 + 2n^1/3 - 15 = 0

solve equation over the set of complex numbers.

Answer by Nate(3500)

n^(1/3) = x
n^(1/3)^2 + 2n^(1/3) - 15 = 0
x^2 + 2x - 15 = 0
(x + 5)(x - 3) = 0
(n^(1/3) + 5)(n^(1/3) - 3) = 0
n^(1/3) + 5 = 0 and n^(1/3) - 3 = 0
n^(1/3) = -5 and n^(1/3) = 3
n = -125 and n = 27